in an arithmetic sequence the ratio of first term to second term is 2:3 .what is the ratio of third to fifth term?
Answer by richwmiller(17219)

Lets try a few examples
2,3,4,5,6
2/3=2/3
4/6=2/3
4,6,8,10,12
4/6=2/3
8/12=2/3
looks like the third and fifth are also 2:3
a,a+d,a+2d,a+3d,a+4d
a/(a+d)=2/3
3a=2a+2d
a=2d
(a+2d)/(a+4d)
2a/3a=2/3
So yes it holds true for all arithmetic series.